A plane can fly 180 mph in calm air. Flying with the wind, the plane can fly 600 mi in the same ammount of time it takes to fly 480 mi against the wind. Find the rate of the wind.
Answer by lwsshak3(11628)

A plane can fly 180 mph in calm air. Flying with the wind, the plane can fly 600 mi in the same ammount of time it takes to fly 480 mi against the wind. Find the rate of the wind.
**
let c=rate of wind
180+c=speed of plane with wind
180-c=speed of plane against wind
travel time=distance/speed
..
600/(180+c)=480/(180-c)
600(180)-600c=480(180)+480c
600(180)-480(180)=600c+480c
180(600-480)=1080c
21600=1080c
c=20 mph
rate of wind=20 mph